The body temperatures of a group of healthy adults have a bell-shaped distribution with a mean of 98.33°F and a standard deviation of 0.65°F. Using the empirical rule, find each approximate percentage below. 

a. What is the approximate percentage of healthy adults with body temperatures within 3 standard deviations of the mean, or between 96.38°F and 100.28°F? 

b. What is the approximate percentage of healthy adults with body temperatures between 97.68 °F and 98.98 °F? 


a. Approximately _______ % of healthy adults in this group have body temperatures within 3 standard deviations of the mean, or between 96.38°F and 100.28 °F. (Type an integer or a decimal. Do not round.)
Answer #1

Solution

Mean = 98.33

standard deviation = 0.65

the Distribution is bell shaped

using empirical rule

(a) between 3 standard deviation of mean approximately 99.7% values falls hence

Answer : Approximately 99.7 % of healthy adults ....

(b) 98.33-97.68= 0.65 and 100.28 - 98.33 = 1.95

hence (34)% + (99.7/2)% = 83.85%

Answer : 83.85%